Solve and check
10(x-7)-3x=x+2

Answer by elima(1433)

10(x-7)-3x=x+2
10x-70-3x=x+2
7x-70=x+2
7x-x=2+70
6x=72
x=12
check;
10((12)-7)-3(12)=(12)+2
120-70-36=14
50-36=14
14=14
:)